OpenStudy (anonymous):

(20)/(x-10)-(3)/(10-x) reduce if possible, assume that no denominator is 0. I need the steps please, thank you.

OpenStudy (anonymous):

Distribute the 20 and the -3 and simplify

OpenStudy (asnaseer):

you should spot that the denominator of the 2nd fraction is just the negative of the denominator of the 1st fraction so that this can be re-arranged as follows:\[\frac{20}{x-10}-\frac{3}{10-x}=\frac{20}{x-10}-\frac{3}{-(x-10)}\]\[=\frac{20}{x-10}+\frac{3}{x-10}\]you should then be able to solve this fairly simply...

OpenStudy (anonymous):

^^ thats easier lol

OpenStudy (asnaseer):

it helps to gaze at a problem for a bit before diving into it.

OpenStudy (asnaseer):

@Cierra-Nicole - do you understand how to complete this problem now?

OpenStudy (anonymous):

yes and no - i have a another problem, i need the complete breakdown.?

OpenStudy (anonymous):